OpenStudy (anonymous):

What is the value of 12C5 ? A. 120 B. 60 C. 95,040 D. 792

OpenStudy (anonymous):

What is the value for c?

OpenStudy (triciaal):

12!/5!

OpenStudy (welshfella):

12*11*10*9*8 ------------ 5*4*3*2*1

OpenStudy (anonymous):

thats not a option @triciaal quit using google to try to find the answer and work it out

OpenStudy (anonymous):

^

OpenStudy (triciaal):

i did not use google

OpenStudy (perl):

$$ 12 C 5 = \frac{12!}{(12-5)! ~5!} $$

OpenStudy (anonymous):

funny cause i just looked it up and it just so happened to give me the "answer" that you gave

OpenStudy (triciaal):

it is just what the notation means from what i remember

OpenStudy (perl):

$$ \Large nCr = \frac{n!}{(n-r)! ~r!} $$

OpenStudy (welshfella):

no i think that the notation for 12P5 - permutations not combinations

OpenStudy (triciaal):

i was trying to remember if i need the 7! like @ perl indicated

OpenStudy (perl):

$$ \Large nPr = \frac{n!}{(n-r)! } $$

OpenStudy (welshfella):

oh right

OpenStudy (perl):

hopefully i remembered correctly :)

OpenStudy (welshfella):

the method i used is a simplification of n! --- (n-r)! r!

OpenStudy (triciaal):

I am not perfect I do make mistakes but I don't appreciate being accused of using google

OpenStudy (welshfella):

yes

OpenStudy (perl):

This is a learning website, mistakes happen ( I make them ). No need for accusations.

OpenStudy (anonymous):

thanks for helping

OpenStudy (welshfella):

yw